Site navigation

The probabilities that A, B, and D can solve a problem independently are 1/3, 1/3, and 1/4 respectively. The probability that only two of them are able to solve the problem is:

Question and options

The probabilities that A, B, and D can solve a problem independently are 1/3, 1/3, and 1/4 respectively. The probability that only two of them are able to solve the problem is:

  1. 536\frac{5}{36}
  2. 736\frac{7}{36}
  3. 1336\frac{13}{36}
  4. 12\frac12

Answer and explanation

Correct answer: B - \(\frac{7}{36}\)
Exactly two: (1/3)(1/3)(3/4) + (1/3)(2/3)(1/4) + (2/3)(1/3)(1/4) = 3/36 + 2/36 + 2/36 = 7/36.

Practice for these exams

  • JIPMAT

Related question papers

Related exam syllabi

More solved explanations in Arithmetic